OpenStudy (anonymous):

what is the inverse relation for 2x+3y=10

OpenStudy (anonymous):

First solve for y: 3y = -2x +10 or y = (-2/3)x + 10/3 Do you know the next step?

OpenStudy (anonymous):

i have to resolve the x and y ?

OpenStudy (mertsj):

2y+3x=10

OpenStudy (anonymous):

but the steps?

OpenStudy (anonymous):

i got what meverett is saying, so what do i do next ?

OpenStudy (anonymous):

2x+3y=10 Subtract 2x from both sides 3y = -2x +10 Divide by 2 y = (-2/3)x + 10

OpenStudy (anonymous):

so that would be its inverse?

OpenStudy (anonymous):

no .... First solve for y: 3y = -2x +10 or y = (-2/3)x + 10/3 Do you know the next step?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i prefer the 3y=-2x+10 since it has no fraction

OpenStudy (anonymous):

okay, take the 3y=-2x+10 and switch the x and y 3X = -2Y + 10 and solve for this new Y .... mom is calling ..... gotta go ... sorry ...

OpenStudy (anonymous):

ok thank you very much

OpenStudy (anonymous):

so the final result would be ?

OpenStudy (mertsj):

Just interchange the x and y and solve for y.